# note-name = , use-sort-key = false

### v3

@@ III.3 Wigner negativity
$\mathcal{V}[W(x,p)]=\frac{1}{2}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}|W(x,p)|\ \mathrm{d}x\ \mathrm{d}p-\frac{1}{2}$ (20)
@@
Quantum decoherence imposed by the nonconservative processes causes the system to lose its quantumness. This translates to the Wigner function losing its negative regions, resulting in a nonnegative $W^{\mathrm{(ss)}}$ as shown in Fig. 2. The decoherence imposed by the one-quantum processes is embodied in the terms $\partial_{x}^{2}W+\partial_{p}^{2}W$ in Eq. (15). These terms cause the negative volume to monotonically decrease, as we can see by noting that their values near a minimum are positive. Ref. [79], for example, has shown this behavior.
@@
Figure 5: Early evolution of the negative volume $\mathcal{V}$ of the Wigner function given by Eq. (20) of some selected states $\left|{\psi}\right\rangle$, with $(\kappa_{1},\gamma_{1},\gamma_{2})=(0.01,0.009,1)$ and $\mathcal{N}$ some normalization constant. The inset shows the evolution of the cat states’ $\mathcal{V}$ for the first hundredth of the simple harmonic period. For the other three states, the evolution of $\mathcal{V}$ over the same time window is simply a monotonous decrease and is not shown. The horizontal axis of the main plot is log-scaled, while that of the inset is linear. A time step of $2002\pi/$ 200 and $100002\pi/$ 10000 are used for the main plot and the inset, respectively. More delicate shapes may be obtained with smaller time steps.
@@
Figure 5 shows some examples of $\mathcal{V}$ increasing due to the $\gamma_{2}$ process in the small $\mathcal{A}$ regime for some initial states showing Wigner negativity. Out of the five states considered, two do not show any increase in $\mathcal{V}$, namely the Fock state $\left|{n=2}\right\rangle$ and the Fock-state superposition $\mathcal{N}\left(\left|{n=2}\right\rangle+\left|{n=5}\right\rangle\right)$. While these do not show $\mathcal{V}$ increasing, $\left|{n=5}\right\rangle$ does, suggesting that the Wigner function of $\left|{n=2}\right\rangle$ smoothes that of $\left|{n=5}\right\rangle$. On the other hand, the two cat states show the feature very early in the evolution, suggesting that their sharp interference pattern in the Wigner function is quickly spread out by the dynamics. The different time it takes for $\mathcal{V}$ to vanish depends on the initial states and how the evolution plays out—here we solely aim to show that it is possible for $\mathcal{V}$ to increase given the right conditions.
IV Evolution Speed and Speedy Parameters
Some interesting features of a self-sustained oscillator are its ability to restore its rhythm after being perturbed and synchronize when coupled. This makes it interesting to study the evolution speed of the system. The system becomes more robust to external perturbations by reaching the limit cycle faster. The attraction speed may further provide some insight into synchronization enhancement, e.g. in speed and robustness. Our system’s dynamics is determined by the nonconservative parameters, as well as the initial state. In this section, we study how these affect how fast the system can reach the limit cycle.
IV.1 The Lindbladian spectral decomposition of the density matrix
$\mathcal{L}^{\dagger}\rho^{\mathrm{(L)}}_{j}=\lambda_{j}^{*}\rho^{\mathrm{(L)}}_{j}\qquad\mathcal{L}\rho^{\mathrm{(R)}}_{j}=\lambda_{j}\rho^{\mathrm{(R)}}_{j}$ (21)
where $\mathrm{Re}\left(\lambda_{j}\right)\leq 0$. Let $\lambda_{j}$ be sorted in order of ascending magnitudes. If a steady state exists, then $\lambda_{0}=0$ and $\rho^{\mathrm{(R)}}_{0}=\rho^{\mathrm{(ss)}}$ since $\mathcal{L}\rho^{\mathrm{(ss)}}=0$. The density matrix at any given time can be expressed in terms of the eigenvalues and eigenmatrices of $\mathcal{L}$ as
$\rho(t)=\rho^{\mathrm{(ss)}}+\sum_{j=1}^{\infty}c_{j}e^{\lambda_{j}t}\rho^{\mathrm{(R)}}_{j}$ (22)
$c_{j}=\frac{\left\langle{\rho^{\mathrm{(L)}}_{j}},{\rho^{\mathrm{(0)}}}\right\rangle_{\mathrm{HS}}}{\left\langle{\rho^{\mathrm{(L)}}_{j}},{\rho^{\mathrm{(R)}}_{j}}\right\rangle_{\mathrm{HS}}}$ (23)
where $\left\langle{\hat{A}},{\hat{B}}\right\rangle_{\mathrm{HS}}=\mathrm{tr}\left({\hat{A}^{\dagger}\hat{B}}\right)$ is the Hilbert-Schmidt norm and $\rho^{\mathrm{(0)}}$ is the initial state. As the system evolves, the real part of $\lambda_{j}$ causes the summation in Eq. (22) to decay toward zero, with the $j=1$ term decaying the slowest. For this reason, $\Delta=\left|\mathrm{Re}\left(\lambda_{1}\right)\right|$ is called the Liouvillian gap or asymptotic decay rate.
@@ IV.1 The Lindbladian spectral decomposition of the density matrix
¯¯$\begin{split}\bar{\bar{{\mathcal{L}}}}&=-i\left(\hat{a}^{\dagger}\hat{a}\otimes\mathbb{1}-\mathbb{1}\otimes\hat{a}\hat{a}^{\dagger}\right)\\ &\qquad+\kappa_{1}\left(\hat{a}^{\dagger}\otimes\hat{a}^{\dagger}-\frac{1}{2}\hat{a}\hat{a}^{\dagger}\otimes\mathbb{1}-\frac{1}{2}\mathbb{1}\otimes\hat{a}\hat{a}^{\dagger}\right)\\ &\qquad+\gamma_{1}\left(\hat{a}\otimes\hat{a}-\frac{1}{2}\hat{a}^{\dagger}\hat{a}\otimes\mathbb{1}-\frac{1}{2}\mathbb{1}\otimes\hat{a}^{\dagger}\hat{a}\right)\\ &\qquad+\gamma_{2}\left(\hat{a}^{2}\otimes\hat{a}^{2}-\frac{1}{2}\hat{a}^{\dagger{2}}\hat{a}^{2}\otimes\mathbb{1}-\frac{1}{2}\mathbb{1}\otimes\hat{a}^{\dagger{2}}\hat{a}^{2}\right)\end{split}$ (24)
@@
¯¯$\begin{split}\bar{\bar{{\mathcal{L}}}}_{(a,b);(n,m)}&=\left\langle{a,b}\middle|{\bar{\bar{{\mathcal{L}}}}}\middle|{m,n}\right\rangle\\ &=\Big{[}-i(m-n)-\frac{\kappa_{1}}{2}(m+n+2)-\frac{\gamma_{1}}{2}(m+n)\\ &\qquad\qquad-\frac{\gamma_{2}}{2}\left(m(m-1)+n(n-1)\right)\Big{]}\delta_{a,m}\delta_{b,n}\\ &\qquad+\kappa_{1}\sqrt{m+1}\sqrt{n+1}\delta_{a,m+1}\delta_{n,m+1}\\ &\qquad+\gamma_{1}\sqrt{m}\sqrt{n}\delta_{a,m-1}\delta_{b,n-1}\\ &\qquad+\gamma_{2}\sqrt{m(m-1)}\sqrt{n(n-1)}\delta_{a,m-2}\delta_{b,n-2}\end{split}$ (25)
@@
where $\delta_{m,n}$ is the Kronecker delta. The vectorized form of $\mathcal{L}$ is neither diagonal nor triangular, so the eigenvalues cannot be straightforwardly extracted and we resort to numerical methods.
While truncating the Hilbert space dimension $N$ means we cannot capture the general dynamics, the results are still valid when the dynamics only involves the first $N$ levels. In Fig. 6 (a), we tile-plot the highest level $n_{\mathrm{hi}}$ occupied by the system’s steady state. Our criterion for this is the lowest level for which $P_{n}^{\mathrm{(ss)}}\leq 10^{-6}$ for all $n_{\mathrm{hi}}<n\leq N$. The value of $n_{\mathrm{hi}}$ depends only on $(\mathcal{A},\mathcal{B})$ and does not depend on the choice of basis parameters. Furthermore, the system’s energy monotonously evolves toward $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$, as can be seen in, for example, Fig. 3. These behaviors make $n_{\mathrm{hi}}$ useful for choosing $N$ to compute the eigenvalues with. We can choose $N$ some levels above $n_{\mathrm{hi}}$ if the system’s dynamics starts at some energy below $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$. Otherwise, we may choose $N$ some levels above the highest occupied level of the initial state. Looking at this another way, given the working regime point, say $(\mathcal{A},\mathcal{B})$, we need at least an $n_{\mathrm{hi}}$-level approximation for the result to describe the dynamics toward the steady state properly, albeit only for limited choices of initial states.
@@
Figure 6: Numerical evaluation of the eigenvalues of the system’s Lindbladian in different working regimes. (a): Tile plot of the highest level $n_{\mathrm{hi}}$ occupied by the steady state $\rho^{\mathrm{(ss)}}$. (b)–(c): Tile plots of the Liouvillian gap $\Delta=\left|\mathrm{Re}\left(\lambda_{1}\right)\right|$ in different working regimes for different truncated Hilbert space dimension $N$. The shaded regions show invalid results: no dynamics from any initial state to the steady state is contained in the first $N$ levels given working regimes points $(\mathcal{A}=\kappa_{1}/\gamma_{2},\mathcal{B}=(\kappa_{1}-\gamma_{1})/\gamma_{2})$ in the shaded regions. (d)–(e): Plots of $\Delta$ against $\mathcal{A}$ for $\mathcal{B}\approx 0.03$ and against $\mathcal{B}$ for $\mathcal{A}\approx 33.11$. The top $y$-axes show $n_{\mathrm{hi}}$ for the given $(\mathcal{A},\mathcal{B})$—the result for a given $N$ is invalid for $x$-axis values for which $n_{\mathrm{hi}}>N$. (f): Scatter plot of the Lindbladian’s eigenvalues $\lambda_{j};j=0,1,\dots,9$ in the complex plane for different $\mathcal{B}$ with $\mathcal{C}=\mathcal{A}/\mathcal{B}=1$. The opaque markers correspond to $N=100$, showing a fully valid result since the highest $n_{\mathrm{hi}}$ is $94$ over the interval. Meanwhile, the faded ones correspond to $N=20$, showing the invalid result most pronounced for $\mathcal{B}=100$ (at the center of the plot). The $x$-axis is linear between 00 and $-0.01$, and logarithmic elsewhere. The basis parameter used for the calculations is $\kappa_{1}=0.1$. Scaling $\kappa_{1}$ by some constant leaves tile plot (a) and $\mathrm{Im}\left(\lambda_{j}\right)$ unchanged, and scales $\mathrm{Re}\left(\lambda_{j}\right)$ by the same constant.
@@
Figures 6 (b)–(c) are tile-plots of $\Delta$ in different working regime points and different $N$. In the shaded regions, $n_{\mathrm{hi}}$ are larger than $N$. As such, the results are invalid there. By increasing $N$, we reveal more valid regions where the values differ from those obtained with a smaller $N$. We can see from the tile plots that a larger $\mathcal{A}$ leads to a smaller $\Delta$. Other features are less obvious and are easier to see by taking a slice of the tile plots. We take a closer look at the dependence on $\mathcal{A}$ by taking the $\mathcal{B}\approx 0.03$ slice (6th column from the left) and plotting the values in Fig. 6 (d). For the values plotted, $N=50$ gives the valid result while $N=20$ gives the invalid result. It turns out that in the weakly nonlinear regime (small $\mathcal{A}$), we can see a small increase in $\Delta$. This increase is then followed by a steady decrease as $\mathcal{A}$ becomes larger. Meanwhile, for a given $\mathcal{A}$ we see a plateau-like dependence on $\mathcal{B}$ which ends at some point as $\mathcal{B}$ approaches $\mathcal{A}$, as shown in Fig. 6 (e) for $\mathcal{A}\approx 33.11$.
@@
As for the validity for a given $N$, contrary to what we described above, the results are pretty similar when $N$ is only a few levels below $n_{\mathrm{hi}}$ where we find the relative error to be as small as 1.7%. Nevertheless, as $n_{\mathrm{hi}}$ becomes significantly larger than $N$, we see a significant deviation of the invalid values. Had we ignored the $n_{\mathrm{hi}}$ condition, we would have erroneously concluded that the Liouvillian gap can increase as $\gamma_{1}$ becomes sufficiently smaller than $\kappa_{1}$. By increasing $N$, the turning point of $\Delta$ decreases and its position shifts toward larger $\mathcal{B}$. As such, the true dynamics represented by $N=\infty$ is that there is no such turning point.
@@
Shifting toward a more general viewpoint, in Fig. 6 (f), we scatter-plot the distribution of $\lambda_{j};j=0,1,\dots,9$ in the complex plane for different $\mathcal{B}$ with $\mathcal{C}=1$. The opaque and faded markers correspond to $N=100$ and $N=20$, respectively. The highest $n_{\mathrm{hi}}$ is $94$ in the plotted interval, so the result for $N=100$ is fully valid. Similar to Fig. 6 (e), a large deviation between the valid and invalid results can be seen when $\mathcal{B}$ becomes large compared to $\mathcal{A}$. The Liouvillian gap $\Delta$ corresponds to a complex conjugate pair of eigenvalues: $\lambda_{1}$ and $\lambda_{2}$. Eigenvalues with larger real-part magnitudes are also multiply real-part degenerate. As $\mathcal{B}$ (and $\mathcal{A}$) increases, the eigenvalues decrease in real-part magnitudes at different rates, so real-part degeneracies higher than 2 are progressively lost. At $\mathcal{B}=100$ and $\mathcal{C}=1$, the first eight closest-to-zero eigenvalues make up five distinct real-part magnitudes, each doubly degenerate. Unlike $n_{\mathrm{hi}}$ and the ratio $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}/\left|\alpha_{\mathrm{c,lc}}\right|^{2}$ shown in Fig. 1, the real parts of the eigenvalues are scaled along the basis parameter. Meanwhile, the imaginary parts are left unchanged, which is consistent with the fact that the nonconservative processes do not affect the coherent oscillation. This can be seen, for example, in Fig. 3 which shows the Wigner function rotating in the phase plane at the simple harmonic frequency.
IV.2 Steady-state time for selected initial states

### v5

@@ III.3 Wigner negativity
$\mathcal{V}[W(x,p)]=\frac{1}{2}\left(\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}|W(x,p)|\ \mathrm{d}x\ \mathrm{d}p-1\right).$ (20)
@@
Quantum decoherence imposed by the nonconservative processes causes the system to lose its quantumness. This translates to the Wigner function losing its negative regions, resulting in a nonnegative $W^{\mathrm{(ss)}}$ as shown in Fig. 2. The decoherence imposed by the one-quantum processes is embodied in the terms $\left(\partial_{x}^{2}+\partial_{p}^{2}\right)W$ in Eq. (15). These terms cause the negative volume to monotonically decrease, as we can see by noting that their values near a minimum are positive. Ref. [81], for example, has shown this behavior.
@@
Figure 5: Early evolution of the negative volume $\mathcal{V}$ of the Wigner function given by Eq. (20) of some selected states $\left|{\psi}\right\rangle$, with $(\kappa_{1},\gamma_{1},\gamma_{2})=(0.01,0.009,1)$ and $\mathcal{N}$ some normalization constant. The inset shows the evolution of the cat states’ $\mathcal{V}$ for the first hundredth of the simple harmonic period. For the other three states, the evolution of $\mathcal{V}$ over the same time window is simply a monotonous decrease and is not shown. The horizontal axis of the main plot is log-scaled, while that of the inset is linear. A time step of $2\pi/200$ and $2\pi/10000$ is used for the main plot and the inset, respectively. More delicate shapes may be obtained with smaller time steps.
On the other hand, the $\gamma_{2}$ process may decrease $W$ and consequently increase $\mathcal{V}$ due to the third-order derivatives in Eq. (15). We can see this by making the KM coefficients precede the differential operators  [^4] As an aside, we can see that the third row is always zero at the phase space origin, meaning that the value of $W$ there never decreases for the system.. For $\mathcal{V}$ to increase, we need $\gamma_{2}$ to be large enough compared to $(\kappa_{1}+\gamma_{1})$ so that the third-term decoherence is larger than the second-term decoherence. As such, it is less likely to observe $\mathcal{V}$ increase in the large $\mathcal{A}$ or $\mathcal{B}$ regime, such as the classical-eligible regime. Furthermore, we need the third-order derivative values to be negative and have larger magnitudes than the second-order derivative values, limiting which states can exhibit increasing $\mathcal{V}$ to those whose Wigner function’s negative regions have quickly decreasing curvatures. Since the terms with second-order derivatives cause $W$ to become less steep with time, it is impossible for $W$ to keep its quickly decreasing curvature. As a result, $\mathcal{V}$ eventually stops being able to increase, and we ultimately have a nonnegative $W$.
Figure 5 shows some examples of $\mathcal{V}$ increasing due to the $\gamma_{2}$ process in the small $\mathcal{A}$ regime for some initial states showing Wigner negativity. Out of the five states considered, two do not show any increase in $\mathcal{V}$, namely the Fock state $\left|{n=2}\right\rangle$ and the Fock-state superposition $\mathcal{N}\left(\left|{n=2}\right\rangle+\left|{n=5}\right\rangle\right)$, where $\mathcal{N}$ is the normalization constant. While these do not show $\mathcal{V}$ increasing, $\left|{n=5}\right\rangle$ does, suggesting that the Wigner function of $\left|{n=2}\right\rangle$ smoothes that of $\left|{n=5}\right\rangle$. On the other hand, the two cat states show the feature very early in the evolution, suggesting that their sharp interference pattern in the Wigner function is quickly spread out by the dynamics. The different time it takes for $\mathcal{V}$ to vanish depends on the initial states and how the evolution plays out. Here, we solely aim to show that it is possible for $\mathcal{V}$ to increase given the right conditions.
IV Evolution Speed and fast convergence regimes
Some interesting features of a self-sustained oscillator are its ability to restore its rhythm after being perturbed and synchronize when coupled. This makes it interesting to study the system’s evolution speed. The system becomes more robust to external perturbations by reaching the limit cycle faster. The attraction speed may also provide insight into synchronization enhancement, e.g., in speed and robustness. Our system’s dynamics is determined by the nonconservative parameters, as well as the initial state. In this section, we investigate how these affect the time it takes the system to reach the limit cycle.
IV.1 The Lindbladian spectra
$\mathcal{L}^{\dagger}\rho^{\mathrm{(L)}}_{j}$ $=\lambda_{j}^{*}\rho^{\mathrm{(L)}}_{j}$, (21a)
$\mathcal{L}\rho^{\mathrm{(R)}}_{j}$ $=\lambda_{j}\rho^{\mathrm{(R)}}_{j}$, (21b)
where $\mathrm{Re}\left(\lambda_{j}\right)\leq 0$, and let $\lambda_{j}$ be sorted in order of ascending magnitudes. If a steady state exists, then $\lambda_{0}=0$ and $\rho^{\mathrm{(R)}}_{0}=\rho^{\mathrm{(ss)}}$ since $\mathcal{L}\rho^{\mathrm{(ss)}}=0$. The density matrix at any given time can be expressed in terms of the eigenvalues and eigenmatrices of $\mathcal{L}$ as
$\rho(t)=\rho^{\mathrm{(ss)}}+\sum_{j=1}^{\infty}c_{j}e^{\lambda_{j}t}\rho^{\mathrm{(R)}}_{j}$, (22)
$c_{j}=\frac{\left\langle{\rho^{\mathrm{(L)}}_{j}},{\rho^{\mathrm{(0)}}}\right\rangle_{\mathrm{HS}}}{\left\langle{\rho^{\mathrm{(L)}}_{j}},{\rho^{\mathrm{(R)}}_{j}}\right\rangle_{\mathrm{HS}}},$ (23)
@@ IV.1 The Lindbladian spectra
$\begin{split}\bar{\bar{{\mathcal{L}}}}&=-i\left(\hat{a}^{\dagger}\hat{a}\otimes\mathbb{1}-\mathbb{1}\otimes\hat{a}\hat{a}^{\dagger}\right)\\ &\qquad+\kappa_{1}\left(\hat{a}^{\dagger}\otimes\hat{a}^{\dagger}-\frac{1}{2}\hat{a}\hat{a}^{\dagger}\otimes\mathbb{1}-\frac{1}{2}\mathbb{1}\otimes\hat{a}\hat{a}^{\dagger}\right)\\ &\qquad+\gamma_{1}\left(\hat{a}\otimes\hat{a}-\frac{1}{2}\hat{a}^{\dagger}\hat{a}\otimes\mathbb{1}-\frac{1}{2}\mathbb{1}\otimes\hat{a}^{\dagger}\hat{a}\right)\\ &\qquad+\gamma_{2}\left(\hat{a}^{2}\otimes\hat{a}^{2}-\frac{1}{2}\hat{a}^{\dagger{2}}\hat{a}^{2}\otimes\mathbb{1}-\frac{1}{2}\mathbb{1}\otimes\hat{a}^{\dagger{2}}\hat{a}^{2}\right),\end{split}$ (24)
@@
$\begin{split}\bar{\bar{{\mathcal{L}}}}_{(a,b);(n,m)}&=\left\langle{a,b}\middle|{\bar{\bar{{\mathcal{L}}}}}\middle|{m,n}\right\rangle\\ &=\Big[-i(m-n)-\frac{\kappa_{1}}{2}(m+n+2)-\frac{\gamma_{1}}{2}(m+n)\\ &\qquad\qquad-\frac{\gamma_{2}}{2}\left(m(m-1)+n(n-1)\right)\Big]\delta_{a,m}\delta_{b,n}\\ &\qquad+\kappa_{1}\sqrt{m+1}\sqrt{n+1}\delta_{a,m+1}\delta_{n,m+1}\\ &\qquad+\gamma_{1}\sqrt{m}\sqrt{n}\delta_{a,m-1}\delta_{b,n-1}\\ &\qquad+\gamma_{2}\sqrt{m(m-1)}\sqrt{n(n-1)}\delta_{a,m-2}\delta_{b,n-2},\end{split}$ (25)
@@
where $\delta_{m,n}$ is the Kronecker delta. The vectorized form of $\mathcal{L}$ is neither diagonal nor triangular, so the eigenvalues cannot be straightforwardly extracted, and we resort to numerical methods.
@@
Figure 6: Numerical evaluation of the eigenvalues of the system’s Lindbladian in different working regimes. (a): Tile plot of the highest level $n_{\mathrm{hi}}$ occupied by the steady state $\rho^{\mathrm{(ss)}}$. (b)–(c): Tile plots of the Liouvillian gap $\Delta=\left|\mathrm{Re}\left(\lambda_{1}\right)\right|$ in different working regimes for different truncated Hilbert space dimension NN. The shaded regions show invalid results: no dynamics from any initial state to the steady state is contained in the first NN levels given working regimes points $(\mathcal{A}=\kappa_{1}/\gamma_{2},\mathcal{B}=(\kappa_{1}-\gamma_{1})/\gamma_{2})$ in the shaded regions. (d)–(e): Plots of $\Delta$ against $\mathcal{A}$ for $\mathcal{B}\approx 0.03$ and against $\mathcal{B}$ for $\mathcal{A}\approx 33.11$. The top $y$-axes show $n_{\mathrm{hi}}$ for the given $(\mathcal{A},\mathcal{B})$; the result for a given $N$ is invalid for $x$-axis values for which $n_{\mathrm{hi}}>N$. (f): Scatter plot of the Lindbladian’s eigenvalues $\lambda_{j};j=0,1,\dots,9$ in the complex plane for different $\mathcal{B}$ with $\mathcal{C}=\mathcal{A}/\mathcal{B}=1$. The filled markers correspond to $N=100$, showing a fully valid result since the highest $n_{\mathrm{hi}}$ is $94$ over the interval. Meanwhile, the empty ones correspond to $N=20$, showing the invalid result most pronounced for $\mathcal{B}=100$ (at the center of the plot). The $x$-axis is linear between 0 and 0.1-0.1 (shaded region), and logarithmic elsewhere. The basis parameter used for the calculations is $\kappa_{1}=0.1$. Scaling $\kappa_{1}$ by some constant leaves tile plot (a) and $\mathrm{Im}\left(\lambda_{j}\right)$ unchanged, and scales $\mathrm{Re}\left(\lambda_{j}\right)$ by the same constant.
While truncating the Hilbert space dimension $N$ means we cannot capture the general dynamics, the results are still valid when the dynamics only involve the first $N$ levels. In Fig. 6(a), we tile-plot the highest level $n_{\mathrm{hi}}$ occupied by the system’s steady state. Our criterion for this is the lowest level for which $P_{n}^{\mathrm{(ss)}}\leq 10^{-6}$ for all $n_{\mathrm{hi}}<n\leq N$. The value of $n_{\mathrm{hi}}$ depends only on $(\mathcal{A},\mathcal{B})$ and does not depend on the choice of basis parameter. Furthermore, the system’s energy monotonously evolves toward $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$, as can be seen in, for example, Fig. 3. These behaviors make $n_{\mathrm{hi}}$ useful for choosing $N$ to compute the eigenvalues with. We can choose $N$ some levels above $n_{\mathrm{hi}}$ if the system’s dynamics starts at some energy below $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$. Otherwise, we may choose $N$ some levels above the highest occupied level of the initial state. Looking at this another way, given the working regime point, say $(\mathcal{A},\mathcal{B})$, we need at least an $n_{\mathrm{hi}}$-level approximation for the result to describe the dynamics toward the steady state properly, albeit only for limited choices of initial states.
@@
Figures 6(b)–(c) are tile-plots of $\Delta$ in different working regime points and different $N$. In the shaded regions, $n_{\mathrm{hi}}$ is larger than $N$. As such, the results are invalid there. By increasing $N$, we reveal more valid regions where the values differ from those obtained with a smaller $N$. We can see from the tile plots that a larger $\mathcal{A}$ leads to a smaller $\Delta$. Other features are less obvious and are easier to see by taking a slice of the tile plots. We take a closer look at the dependence on $\mathcal{A}$ by taking the $\mathcal{B}\approx 0.03$ slice (6th column from the left) and plotting the values in Fig. 6(d). For the values plotted, $N=50$ gives the valid result while $N=20$ gives the invalid result. It turns out that in the weakly nonlinear regime (small $\mathcal{A}$), we can see a small increase in $\Delta$. This increase is then followed by a steady decrease as $\mathcal{A}$ becomes larger. Meanwhile, for a given $\mathcal{A}$ we see a plateau-like dependence on $\mathcal{B}$ which ends at some point as $\mathcal{B}$ approaches $\mathcal{A}$, as shown in Fig. 6(e) for $\mathcal{A}\approx 33.11$.
@@
As for the validity for a given $N$, contrary to what we described above, the results are pretty similar when $N$ is only a few levels below $n_{\mathrm{hi}}$ where we find the relative error to be as small as 1.7%. Nevertheless, as $n_{\mathrm{hi}}$ becomes significantly larger than $N$, we see a significant deviation of the invalid values. Had we ignored the $n_{\mathrm{hi}}$ condition, we would have erroneously concluded that the Liouvillian gap can increase as $\gamma_{1}$ becomes sufficiently smaller than $\kappa_{1}$, forming a turning point that signifies a phase transition [83]. By increasing $N$, the turning point of $\Delta$ decreases and its position shifts toward larger $\mathcal{B}$. As such, the true dynamics obtained as $N\rightarrow\infty$ is that there is no such turning point.
@@
Shifting toward a more general viewpoint, in Fig. 6(f), we scatter-plot the distribution of $\lambda_{j};j=0,1,\dots,9$ in the complex plane for different $\mathcal{B}$ with $\mathcal{C}=1$. This shows the system’s Liouvillian spectra close to Hopf bifurcation (when $\kappa_{1}=\gamma_{1}$), whose general features have been previously discussed by Ref. [38]. The opaque and faded markers correspond to $N=100$ and $N=20$, respectively. The highest $n_{\mathrm{hi}}$ is $94$ in the plotted interval, so the result for $N=100$ is fully valid. Similar to Fig. 6(e), a large deviation between the valid and invalid results can be seen when $\mathcal{B}$ becomes large compared to $\mathcal{A}$. The Liouvillian gap $\Delta$ corresponds to a complex conjugate pair of eigenvalues: $\lambda_{1}$ and $\lambda_{2}$. Eigenvalues with larger real-part magnitudes are also multiply real-part degenerate. As $\mathcal{B}$ (and $\mathcal{A}$) increases, the eigenvalues decrease in real-part magnitudes at different rates, so real-part degeneracies higher than 2 are progressively lost. At $\mathcal{B}=100$ and $\mathcal{C}=1$, the closest-to-zero eigenvalues become doubly degenerate. They lie close to the imaginary axis of the plot, indicating that the system is close to the classical limit, for which the Liouvillian eigenvalues lie on the imaginary axis [38]. Unlike $n_{\mathrm{hi}}$ and the ratio $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}/\left|\alpha_{\mathrm{c,lc}}\right|^{2}$ shown in Fig. 1, the real parts of the eigenvalues are scaled along the basis parameter. Meanwhile, the imaginary parts are left unchanged, which is consistent with the fact that the nonconservative processes do not affect the oscillation (i.e., the rotation of the system’s Wigner function). This can be seen, for example, in Fig. 3, which shows the Wigner function rotating in the phase plane at the simple harmonic frequency.
IV.2 Steady-state times for selected initial states